OpenStudy (anonymous):

find the approximate equation of the line tangent to the graph of y=xcscx at x=pi/6

OpenStudy (anonymous):

@jdoe0001 if you could help i would appreciate it

OpenStudy (freckles):

You need to differentiate y

OpenStudy (anonymous):

yes i get it to be -cscxtanx+cscx

OpenStudy (freckles):

where this tan come in?

OpenStudy (freckles):

\[y'=(x)'\csc(x)+x(\csc(x))' \\ y'=1 \csc(x)+x(-\csc(x)\cot(x))\]

OpenStudy (anonymous):

yes thats correct and then i plug in pi/6 and end up with -1.4641

OpenStudy (freckles):

now you plug in pi/6 to get the slope of the line

OpenStudy (anonymous):

which is not one of my anwser choices

OpenStudy (freckles):

csc(pi/6)=2 and cot(pi/6)=sqrt(3)

OpenStudy (freckles):

\[y'|_{x=\frac{\pi}{6}}=2-\frac{\pi}{6} (2)(\sqrt{3})=2-\frac{\sqrt{3} \pi}{3}\]

OpenStudy (freckles):

and yes that gives me a totally different slope number than what you got

OpenStudy (freckles):

approximately .1862

OpenStudy (anonymous):

yes thats correct also i realized my mistake thank your for your time and effort as i appreciate your assistance @freckles

OpenStudy (freckles):

np